OpenStudy (anonymous):

Determine the centre (h,k) and the radius r of the circle: 4x^2 – x + 4y^2 – 7y – 9 = 0

myininaya (myininaya):

write in (x-h)^2+(y-k)^2=r^2 form

OpenStudy (anonymous):

that's it?

OpenStudy (anonymous):

not going to divide by 4, complete the square and all that? c'mon myininaya!

myininaya (myininaya):

you want me to do their problem?

OpenStudy (anonymous):

well of course! the form is easy, it is getting there that is hard

OpenStudy (anonymous):

how do i play this song on the piano? hit the right keys...

myininaya (myininaya):

i figure they wanted to try first

OpenStudy (anonymous):

yeah well this one is really ugly. maybe there is a typo?

myininaya (myininaya):

but if you insist that i do it for them \[(4x^2-x)+(4y^2-7y)=9\] \[(4x^2-\frac{4}{4}x)+(4y^2-\frac{4}{4}*7y)=9\] \[4(x^2-\frac{1}{4}x)+4(y^2+\frac{1}{4}*7y)=9\] \[4(x^2-\frac{1}{4}x+(\frac{1}{2*4})^2)+4(y^2+\frac{7}{4}y+(\frac{7}{2*4})^2)=9+4*(\frac{1}{2*4})^2+4*(\frac{7}{2*4})^2\] \[4(x-\frac{1}{8})^2+4(y+\frac{7}{8})^2=9+4*\frac{1}{64}+4*\frac{49}{64}\] \[4(x-\frac{1}{8})^2+4(y+\frac{7}{8})^2=9+\frac{1}{16}+\frac{49}{16}\] \[4(x-\frac{1}{8})^2+4(y+\frac{7}{8})^2=9*\frac{16}{16}+\frac{50}{16}\] \[4(x-\frac{1}{8})^2+4(y+\frac{7}{8})^2=\frac{194}{16}\] \[\frac{1}{4}[4(x-\frac{1}{8})^2+4(y+\frac{7}{8})^2]=\frac{1}{4}*\frac{196}{16}\] \[(x-\frac{1}{8})^2+(y+\frac{7}{8})^2=\frac{49}{16}\]

myininaya (myininaya):

Center is \[(\frac{1}{8},-\frac{7}{8})\] \[r^2=\frac{49}{16}=> r=\frac{7}{4}\]

OpenStudy (anonymous):

Thanks for you help :)

OpenStudy (zarkon):

http://alturl.com/452gu

OpenStudy (anonymous):

that must have kept you busy

myininaya (myininaya):

not really it was easy

myininaya (myininaya):

interesing wikepedia and my answer have a different radius

myininaya (myininaya):

oops i see my center is diffent too i changed something that i wasnt suppose to

OpenStudy (anonymous):

yeah i cheated too at first. that is why i said it was really ugly. but maybe i entered it wrong

myininaya (myininaya):

a minus to a plus

OpenStudy (anonymous):

kept you occupied for a while though, didn't it?

myininaya (myininaya):

see second to third i changed something

myininaya (myininaya):

no it didn't

myininaya (myininaya):

but i got the wrong answer lol

myininaya (myininaya):

like always

OpenStudy (anonymous):

i would have divided by 4 first

myininaya (myininaya):

we could have done that first

myininaya (myininaya):

i'm not typing that over again just to show john my mistake

myininaya (myininaya):

he can read this and he can make the correction himself

OpenStudy (anonymous):

no don't bother

OpenStudy (anonymous):

btw did you notice that the wiz gave ictrees a medal?

OpenStudy (anonymous):

lol

myininaya (myininaya):

why is the radius different though?

myininaya (myininaya):

no lol

myininaya (myininaya):

my radius and wikepedia's radius differ by like .03125

myininaya (myininaya):

(x-\frac{1}{8})^2+(y-\frac{7}{8})^2=\frac{49}{16} i think this is suppose to be it

OpenStudy (anonymous):

it got \[\sqrt{\frac{97}{32}}\]

myininaya (myininaya):

darn

myininaya (myininaya):

\[(x-\frac{1}{8})^2+(y-\frac{7}{8})^2=\frac{49}{16}\]

myininaya (myininaya):

really?

OpenStudy (anonymous):

hmmm

OpenStudy (zarkon):

yes

myininaya (myininaya):

let me look at this again

OpenStudy (anonymous):

complete the square is a pain that is why

OpenStudy (anonymous):

and that is why we all love \[-\frac{b}{2a}\]

myininaya (myininaya):

ok wikepedia is right lol

myininaya (myininaya):

its hard to beat a computer

OpenStudy (anonymous):

you must be tired. especially if you are calling wolfram wiki

myininaya (myininaya):

oh lol

myininaya (myininaya):

yes they are both w words

OpenStudy (anonymous):

i am too. off to bed.

myininaya (myininaya):

i got them confused

myininaya (myininaya):

i like my wrong radius better thought since it had a pretty answer

myininaya (myininaya):

though*

OpenStudy (zarkon):

it was nicer :)

myininaya (myininaya):

why are teachers so mean algebra students can't do this

OpenStudy (zarkon):

i don't know

myininaya (myininaya):

my students would hate me for those fractions

OpenStudy (zarkon):

yep...that's for sure

OpenStudy (anonymous):

i am sure your students think you are cute

myininaya (myininaya):

lol i hope they do because thats what im there for

myininaya (myininaya):

to look cute

myininaya (myininaya):

i remember when i was a student correcting my teachers when they made a mistake it made me feel so smart

myininaya (myininaya):

it gave me a high

myininaya (myininaya):

i hate when i make mistakes

myininaya (myininaya):

zarkon i see a fun problem

OpenStudy (zarkon):

I don't make mistakes...I make it look like I'm making a mistake to see if my students are paying attention ;) (at least that is what I tell them )

OpenStudy (zarkon):

where is that

myininaya (myininaya):

lol